# Hybrid classical-quantum communication networks

### v1

@@ Abstract
Quantum communication Quantum networking Quantum key distribution Fiber optic communication Multiplexing
@@
[1]organization=Elmore Family School of Electrical and Computer Engineering and Purdue Quantum Science and Engineering Institute,city=West Lafayette, state=Indiana, postcode=47907, country=USA
@@ 1 Introduction
The prediction by Charles Kao and George Hockham in 1966 that ultra-low-loss silica fiber could enable long-distance information transmission has had a profound impact on society [116]. During the past five decades, sustained efforts by scientists and engineers around the world have transformed fiber optic communication from simple point-to-point links that span tens of kilometers [221, 24] into an extensive, interconnected global network that now supports most Internet traffic.
@@
More recently, the concept of a quantum Internet has emerged through theoretical proposals and preliminary experiments [120, 245]. The ultimate goal is to connect remote quantum information processing devices by exchanging both quantum and classical information. Harnessing the power of quantum superposition and entanglement, the quantum Internet could enable groundbreaking applications, including distributed quantum computing, quantum-secure communication, and distributed quantum sensing. However, quantum systems are notoriously prone to decoherence due to interactions with the environment. Currently, photons in the optical domain, where black-body radiation is negligible, are considered to be the most promising information carriers for long-distance quantum communication. Consequently, the future quantum Internet is most likely to be built upon fiber optic infrastructures, along with free-space optical links. Thus, it is essential to understand to what extent we can leverage existing classical fiber-optic communication technologies and infrastructure to implement various quantum communication protocols.
@@
The primary objectives of a communication system are to transmit information efficiently, reliably, and securely. Technological advances in this field are largely driven by efforts to address two key challenges: extending communication distances and optimizing network resource allocation among multiple users. The evolution of classical fiber optic communication networks can be categorized into three main eras [14, 252]: (1) Regenerated Direct-Detection Era: This phase extended communication distances using classical regenerators. During this initial period, communication capacity was primarily limited by the bandwidth of transceivers. (2) Amplified Dispersion-Managed Era: Characterized by the widespread adoption of wavelength-division multiplexing (WDM) and practical optical amplifiers, such as erbium-doped fiber amplifiers (EDFAs), this era introduced more cost-effective and scalable solutions. (3) Amplified Coherent Communication Era: Marked by the introduction of coherent optical receivers, this phase enabled longer regenerator spacing and significantly improved spectral efficiency. Looking ahead, while continuing to increase the capacity of single-fiber links may not fully meet the exponential demand for data transmission, parallel spatial paths through space-division multiplexing (SDM) [129] present a promising solution.
@@
Quantum communication, in contrast, remains in its infancy. Researchers are actively exploring the theoretical foundations of a quantum Internet, addressing aspects ranging from the physical layer to network control architecture. However, experimental advances have lagged considerably behind theoretical progress. Even implementing the most basic communication task—transmitting a qubit over a lossy channel deterministically and with high fidelity [102]—poses immense challenges, let alone developing more advanced protocols. To date, the primary exceptions are protocols that compatible with probabilistic quantum communication. A prominent example is quantum key distribution (QKD) [21, 72], which allows remote users to establish a shared cryptographic key using a probabilistic quantum channel in conjunction with an authenticated classical channel. In fact, QKD and related quantum cryptographic protocols are expected to be among the most impactful and practical applications of quantum networks in the near term.
@@
On the one hand, given the limited near-term applications of quantum networking, a viable path toward commercialization involves the integration of quantum communication protocols into existing fiber optic networks. Significant research has been dedicated to this approach, from investigating the coexistence of classical and quantum signals within the same fiber via WDM technology to exploring systems that enable simultaneous classical and quantum communications. On the other hand, innovative technologies developed for a future quantum Internet—such as low-loss fiber-optic devices, optical memory, and highly sensitive detectors—could also bring transformative advancements to classical communication networks. Our vision for the Internet’s future is that of a heterogeneous communication network, thoughtfully designed to seamlessly support both classical and quantum communications.
@@ 2 Fundamentals of classical and quantum networking
Two main challenges in large-scale communication networks are ensuring reliable information transmission over lossy and noisy channels, and optimizing the efficient use of network resources. In this section, we present concise overviews of key technologies that have been developed and proposed in both classical fiber-optic and emerging quantum networking.
@@ 2.1 Fundamentals of classical fiber-optic networks
A major advantage of optical communication is its vast bandwidth capacity. However, early forms of long-distance optical communication were limited in application due to several factors: communication rates were constrained by slow encoding processes, and reliability was compromised by unpredictable weather conditions. Widespread long-distance optical communication only became feasible after Corning scientists Dr. Robert Maurer, Dr. Peter Schultz, and Dr. Donald Keck developed ultra-low-loss silica fiber in 1970, following predictions by Charles Kao and George Hockham in 1966 [116]. Since then, the information capacity of fiber-optic networks—measured by the bit rate–distance product (BL)—has grown exponentially, as illustrated in Fig. 1 (note the acceleration in growth rate around 1977) [11]. In the following section, we introduce the fundamentals of classical fiber-optic communication networks.
Figure 1: Information capacity of communication networks measured in the bit rate–distance product (BL) has grown exponentially during the period of 1840-2015 (Images reproduced with permission from a Creative Commons Attribution 4.0 International License (http://creativecommons.org/licenses/by/4.0/) [11])
@@ 2.1.1 The basic structure of an optical communication system
Figure 2 illustrates the basic structure of a typical optical communication system. On the transmitter side, a source encoder can be employed to compress the input message, enhancing communication efficiency. According to Shannon’s source coding theorem [211], in the asymptotic case, a N-bit message from an independent and identically distributed (i.i.d.) source can be compressed into approximately $NH(X)$ bits with minimal risk of information loss, where $H(X)$ is the source entropy. On the receiver side, a source decoder is used to reconstruct the original message from the compressed data.
@@
In early fiber-optic communication systems, low-cost broadband sources such as light-emitting diodes (LEDs) were commonly used. However, to make more efficient use of the optical bandwidth, modern systems typically employ narrowband sources with a linewidth smaller than the modulation bandwidth. This approach maximizes data-carrying capacity and improves overall system performance.
@@
Direct detection is simpler in structure and easier to implement, making it advantageous in certain applications. However, optical coherent detection is becoming the dominant choice in long-distance fiber-optic communications due to its ability to extend regenerator spacing and enhance spectral efficiency [118]. A key advantage of coherent detection is the intrinsic filtering provided by the LO: only signals that match the LO in polarization, spectral, and spatiotemporal mode contribute significantly to the output electrical signal. This filtering capability is especially valuable in quantum communication over dense wavelength-division-multiplexed fiber networks carrying strong classical traffic [192, 122, 71], as well as in free-space optical links [103]. In those cases, noise photons in other modes can be effectively suppressed in coherent detection.
@@ 2.1.2 Fiber loss management
An effective communication channel ensures error-free transmission of messages from the sender to the receiver. In fiber-optic communication, fiber loss is a key factor influencing the range of transmission. The advent of low-loss optical fibers transformed fiber-optic networks, establishing them as the backbone of the modern Internet. In recognition of this milestone, Charles Kuen Kao was awarded half of the Nobel Prize in Physics in 2009 “for groundbreaking achievements concerning the transmission of light in fibers for optical communication” ($https://www.nobelprize.org/prizes/physics/2009/summary/$). As shown in Fig. 1, the rate of exponential growth in communication capacity significantly accelerated around the 1970s with the global deployment of fiber-optic networks [11].
@@
Even with ultra-low-loss fibers, such as those with an attenuation of 0.14 dB/km, communication over long fiber links becomes impractical over distances of several hundred kilometers. For instance, in a 500 km link, the attenuation reaches approximately 70 dB, meaning that, on average, only one photon out of every ten million would successfully pass through the fiber. To overcome this limitation and extend the reach of optical communications, regenerators and optical amplifiers have been introduced.
@@
Although the concept of optical amplification dates back to the 1960s, its widespread application in fiber-optic communication networks did not begin until the 1990s, following the development of EDFA [158, 64]. Doped with the rare-earth element erbium, EDFAs offer low-noise amplification in the 1530–1565 nm wavelength range, which corresponds to the low-loss C-band of silica-based optical fibers. Today, EDFAs are the most widely deployed fiber amplifiers in commercial systems.
@@ 2.1.3 Fiber dispersion management
Once the distance limitation caused by fiber attenuation was overcome through the introduction of optical amplifiers, chromatic dispersion (CD) emerged as one of the most significant impairments in fiber-optic communication systems [49].
As a light pulse propagates through an optical fiber, chromatic dispersion causes different frequency components of the pulse to travel at slightly different speeds. This results in pulse broadening, which leads to inter-symbol interference (ISI) as adjacent pulses begin to overlap. ISI imposes a limit on the maximum error-free communication distance. The challenge becomes more pronounced at higher transmission rates: as the transmission rate increases, pulse duration shortens, and spectral width broadens, causing the dispersion-limited distance to decrease quadratically with the symbol rate. For instance, the chromatic dispersion of standard single-mode fiber (SMF) is approximately 17 ps/nm/km at 1550 nm. When the transmission rate increases from 2.5 Gb/s to 10 Gb/s, the dispersion-limited transmission distance decreases from around 800 km to just 60 km, even with chirp-free light sources [252]. This underscores the importance of dispersion management (DM) in high-speed communication systems.
@@ 2.1.4 Multiuser lightwave networks
Ubiquitous in optical networks for several decades now [109, 34, 26], WDM has historically been divided into “coarse” WDM (CWDM) and “dense WDM” (DWDM). ITU specifies the former as 20 nm-wide channels [236] and the latter as channel spacings ranging from 12.5 GHz to multiples of 100 GHz, with center frequencies equal to integer channel multiples away from 193.1 THz ($\sim$1552.5 nm) [238]. From a practical perspective, the primary difference between CWDM and DWDM networking centers on cost and performance tradeoffs. The much tighter spectral tolerances of DWDM require lasers and filters with GHz-scale accuracies, but in turn can support many more users: for example, the 12.5 GHz DWDM grid corresponds to more than 900 channels in the combined C- and L-bands (1530–1625 nm), compared to a mere five CWDM channels in that same range.
@@
An even more flexible form of DWDM has emerged in recent years—namely “elastic” or “flex-grid” optical networking [85, 110]. Rather than relying on fixed channel center frequencies and widths, flex-grid supports spectral channel of a variety of widths, enabling precise allocation of bandwidth to support disparate user needs; e.g., a high-bandwidth demand can receive a single ultrabroad 1 THz channel, while a low-bandwidth user concurrently obtains a 12.5 GHz one—there is no need to force them onto a fixed grid that may leave critical resources underutilized. ITU recommends central frequencies on a 6.25 GHz grid and slot widths in multiples of 12.5 GHz, but otherwise allows for any set of nonoverlapping channels [238]. In practice, flex-grid implementations are enabled by technology such as wavelength-selective switches (WSSs) that combine Fourier-transform pulse shaping [246, 247] with sub-spot-size gratings for fully programmable control of the amplitude, phase, and output fiber on a frequency-by-frequency basis (limited only by spectral resolution) [195, 159]. With such technology, it is possible to not only realize specific instantiations of flexible DWDM allocations, but also reconfigure slots on demand as bandwidth requirements evolve.
@@
An additional multiplexing approach available to optical networks, time-division multiplexing (TDM) assigns users slots in the time domain, typically interleaved such that $N$ users each communicating at rate $R_{0}$ multiplex into a single line rate of $NR_{0}$. Also enjoying a longstanding history in lightwave communications [225, 208, 206], TDM can be divided into electrical and optical versions, distinguished by whether the (de)multiplexing operations are performed in the electronic or photonic domains, respectively. In the former case, multiple data streams are combined digitally and modulated onto a optical field; at the receiver, the data stream is detected at the full line rate and then demultiplexed into constituent tributary channels digitally. In the optical approach, the tributary channels are modulated onto optical fields separately, then combined, transmitted, and demultiplexed optically before optical-to-electrical (OE) conversion at separate detectors. As an entirely digital technique, electrical TDM requires no modifications to the optical architecture, yet the line rate $NR_{0}$ is then constrained by the bandwidth of OE conversion (either on the modulation or detection front). By contrast, optical TDM requires OE bandwidths only at the tributary rate $R_{0}$, thus in principle enabling much higher total line rates. Nevertheless, all-optical TDM still requires resources at the full bandwidth $NR_{0}$, including short pulses, delay lines, and fast switches. Such technological difficulties combined with the continued increase in digital signal processing bandwidths have prevented optical TDM from making a significant commercial impact.
@@
Finally, moving beyond standard SMF to novel fiber types is opening up additional multiuser opportunities in SDM, where users are assigned distinct transverse spatial modes in, e.g., the same core of a multimode fiber (MMF) or separate cores in the common cladding of a multicore fiber [196, 130, 182]. While the cost of system upgrades makes the level of commercial SDM adoption unclear at the moment, SDM’s ability to multiply the transmission capacity of individual cables by potentially orders of magnitude suggest a compelling niche in space-constrained and high-bandwidth links.

### v2

@@ Abstract
Quantum communication Quantum networking Quantum key distribution Fiber-optic communication Multiplexing
@@
[1]organization=Elmore Family School of Electrical and Computer Engineering and Purdue Quantum Science and Engineering Institute, Purdue University,city=West Lafayette, state=Indiana, postcode=47907, country=USA
@@ 1 Introduction
The prediction by Charles Kao and George Hockham in 1966 that ultralow-loss silica fiber could enable long-distance information transmission has had a profound impact on society [1]. During the past five decades, sustained efforts by scientists and engineers around the world have transformed fiber-optic communication from simple point-to-point links that span tens of kilometers [2, 3] into an extensive, interconnected global network that now supports most Internet traffic.
@@
More recently, the concept of a quantum Internet has emerged through theoretical proposals and preliminary experiments [4, 5]. The ultimate goal is to connect remote quantum information processing devices by exchanging both quantum and classical information. Harnessing the power of quantum superposition and entanglement, the quantum Internet could enable groundbreaking applications, including distributed quantum computing, quantum-secure communication, and distributed quantum sensing. However, quantum systems are notoriously prone to decoherence due to interactions with the environment. Currently, photons in the optical domain, where blackbody radiation is negligible, are considered to be the most promising information carriers for long-distance quantum communication. Consequently, the future quantum Internet is most likely to be built upon fiber-optic infrastructures, along with free-space optical links. Thus, it is essential to understand to what extent we can leverage existing classical fiber-optic communication technologies and infrastructure to implement various quantum communication protocols.
@@
The primary objectives of a communication system are to transmit information efficiently, reliably, and securely. Technological advances in this field are largely driven by efforts to address two key challenges: extending communication distances and optimizing network resource allocation among multiple users. The evolution of classical fiber-optic communication networks can be categorized into three main eras [6, 7]. (1) Regenerated Direct-Detection Era. This phase extended communication distances using classical regenerators. During this initial period, communication capacity was primarily limited by the bandwidth of transceivers. (2) Amplified Dispersion-Managed Era. Characterized by the widespread adoption of wavelength-division multiplexing (WDM) and practical optical amplifiers, such as erbium-doped fiber amplifiers (EDFAs), this era introduced more cost-effective and scalable solutions. (3) Amplified Coherent Communication Era. Marked by the introduction of coherent optical receivers, this phase enabled longer regenerator spacing and significantly improved spectral efficiency. Looking ahead, while continuing to increase the capacity of single-fiber links may not fully meet the exponential demand for data transmission, parallel spatial paths through space-division multiplexing (SDM) [8] present a promising solution.
@@
Quantum communication, in contrast, remains in its infancy. Researchers are actively exploring the theoretical foundations of a quantum Internet, addressing aspects ranging from the physical layer to network control architecture. However, experimental advances have lagged considerably behind theoretical progress. Even implementing the most basic communication task—transmitting a qubit over a lossy channel deterministically and with high fidelity [9]—poses immense challenges, let alone developing more advanced protocols. To date, the primary exceptions are protocols compatible with probabilistic quantum communication. A prominent example is quantum key distribution (QKD) [10, 11], which allows remote users to establish a shared cryptographic key using a probabilistic quantum channel in conjunction with an authenticated classical channel. In fact, QKD and related quantum cryptographic protocols are expected to be among the most impactful and practical applications of quantum networks in the near term.
@@
On the one hand, given the limited near-term applications of quantum networking, a viable path toward commercialization involves the integration of quantum communication protocols into existing fiber-optic networks. Significant research has been dedicated to this approach, from investigating the coexistence of classical and quantum signals within the same fiber via WDM technology to exploring systems that enable simultaneous classical and quantum communications. On the other hand, innovative technologies developed for a future quantum Internet—such as low-loss fiber-optic devices, optical memory, highly sensitive detectors, and distributed entanglement—could also bring transformative advancements to classical communication networks. Our vision for the Internet’s future is that of a heterogeneous communication network, thoughtfully designed to seamlessly support both classical and quantum communications.
@@ 2 Fundamentals of classical and quantum networking
Two main challenges in large-scale communication networks are ensuring reliable information transmission over lossy and noisy channels, and optimizing the efficient use of network resources. In this section, we present concise overviews of key technologies that have been developed and proposed in both classical fiber-optic and emerging quantum networks.
@@ 2.1 Fundamentals of classical fiber-optic networks
A major advantage of optical communication is its vast bandwidth capacity. However, early forms of long-distance optical communication were limited in application due to several factors: communication rates were constrained by slow encoding processes, and reliability was compromised by unpredictable weather conditions. Widespread long-distance optical communication only became feasible after Corning scientists Robert Maurer, Peter Schultz, and Donald Keck developed ultralow-loss silica fiber in 1970, following predictions by Charles Kao and George Hockham in 1966 [1]. Since then, the information capacity of fiber-optic networks—measured by the bit rate–distance product (BL)—has grown exponentially, as illustrated in Fig. 1 (note the acceleration in growth rate around 1977) [14]. In the following section, we introduce the fundamentals of classical fiber-optic communication networks.
Figure 1: Information capacity of communication networks measured in the bit rate–distance product (BL) has grown exponentially during the period of 1840–2015. (Image reproduced with permission from a Creative Commons Attribution 4.0 International License (http://creativecommons.org/licenses/by/4.0/) [14].)
@@ 2.1.1 The basic structure of an optical communication system
Figure 2 illustrates the basic structure of a typical optical communication system. On the transmitter side, a source encoder can be employed to compress the input message, enhancing communication efficiency. According to Shannon’s source coding theorem [15], in the asymptotic case, an $N$-bit message from an independent and identically distributed (i.i.d.) source can be compressed into approximately $NH(X)$ bits with minimal risk of information loss, where $H(X)$ is the source entropy. On the receiver side, a source decoder is used to reconstruct the original message from the compressed data.
@@
In early fiber-optic communication systems, broadband sources such as light-emitting diodes (LEDs) were commonly used due to their simple driver requirements and low cost [16]. However, to make more efficient use of the optical bandwidth, in modern coherent WDM systems, narrowband sources with a linewidth smaller than the modulation bandwidth are typically employed [17]. This approach maximizes data-carrying capacity and improves overall system performance. Despite this evolution, LEDs remain relevant for short-distance links and are anticipated to play a key role in emerging technologies such as the Internet of Things (IoT) and 6G wireless networks [18].
@@
Direct detection is simpler in structure and easier to implement, making it advantageous in certain applications. However, optical coherent detection is becoming the dominant choice in long-distance fiber-optic communications due to its ability to extend regenerator spacing and enhance spectral efficiency [19]. A key advantage of coherent detection is the intrinsic filtering provided by the LO: only signals that match the LO in polarization, spectral, and spatiotemporal mode contribute significantly to the output electrical signal. This filtering capability is especially valuable in quantum communication over dense WDM fiber networks carrying strong classical traffic [20, 21, 22], as well as in free-space optical links [23]. In those cases, noise photons in other modes can be effectively suppressed in coherent detection.
@@ 2.1.2 Fiber loss management
An effective communication channel ensures error-free transmission of messages from the sender to the receiver. In fiber-optic communication, fiber loss is a key factor influencing the range of transmission. The advent of low-loss optical fibers transformed fiber-optic networks, establishing them as the backbone of the modern Internet. In recognition of this milestone, Charles Kuen Kao was awarded half of the Nobel Prize in Physics in 2009 “for groundbreaking achievements concerning the transmission of light in fibers for optical communication” [24]. As shown in Fig. 1, the rate of exponential growth in communication capacity significantly accelerated around the 1970s with the global deployment of fiber-optic networks [14].
@@
Even with ultralow-loss fibers, such as those with an attenuation of 0.14 dB/km, communication over long fiber links becomes impractical over distances of several hundred kilometers. For instance, in a 500 km link, the attenuation reaches approximately 70 dB, meaning that, on average, only one photon out of every ten million would successfully pass through the fiber. To overcome this limitation and extend the reach of optical communications, regenerators and optical amplifiers have been introduced.
@@
Although the concept of optical amplification dates back to the 1960s, its widespread application in fiber-optic communication networks did not begin until the 1990s, following the development of EDFAs [30, 31]. Since then, the EDFA has become the most widely deployed optical amplifier in commercial systems. Doped with the rare-earth element erbium, standard EDFAs offer low-noise amplification in the 1530–1565 nm wavelength range, which corresponds to the low-loss C-band of silica-based optical fibers.
To accommodate growing bandwidth demands, the operational range of EDFAs was later extended to the long-wavelength L-band (1565–1625 nm). The resulting C+L-band configuration significantly increases the number of WDM channels that can be supported a single fiber strand [32]. While efforts to extend amplification beyond the C+L-band have yet to see widespread deployment, such advancements are expected to play a significant role in future capacity upgrades of optical networks [33].
Typically, commercial EDFAs offer gains in the range of 20–30 dB, enabling signal transmission over distances up to approximately 100 km in standard single-mode fiber without regeneration. However, state-of-the-art EDFAs can achieve gains exceeding 40 dB, as demonstrated by commercially available systems [34].
@@ 2.1.3 Fiber dispersion management
Once the distance limitation caused by fiber attenuation was overcome through the introduction of optical amplifiers, chromatic dispersion emerged as one of the most significant impairments in fiber-optic communication systems [36].
As a light pulse propagates through an optical fiber, chromatic dispersion causes different frequency components of the pulse to travel at slightly different speeds. This results in pulse broadening, which leads to intersymbol interference (ISI) as adjacent pulses begin to overlap. ISI imposes a limit on the maximum error-free communication distance. The challenge becomes more pronounced at higher transmission rates: as the transmission rate increases, pulse duration shortens, and spectral width broadens, causing the dispersion-limited distance to decrease quadratically with the symbol rate. For instance, the chromatic dispersion of standard single-mode fiber (SMF) is approximately 17 ps/nm/km at 1550 nm. When the transmission rate increases from 2.5 Gb/s to 10 Gb/s, the dispersion-limited transmission distance decreases from around 800 km to just 60 km, even with chirp-free light sources [7]. This underscores the importance of dispersion management in high-speed communication systems.
@@ 2.1.4 Multiuser lightwave networks
Ubiquitous in optical networks for several decades now [38, 39, 40], WDM has historically been divided into “coarse” WDM (CWDM) and “dense WDM” (DWDM). ITU specifies the former as 20 nm-wide channels [41] and the latter as channel spacings ranging from 12.5 GHz to multiples of 100 GHz, with center frequencies equal to integer channel multiples away from 193.1 THz ($\sim$1552.5 nm) [42]. From a practical perspective, the primary difference between CWDM and DWDM networking centers on cost and performance tradeoffs. The much tighter spectral tolerances of DWDM require lasers and filters with GHz-scale accuracies, but in turn can support many more users: for example, the 12.5 GHz DWDM grid corresponds to more than 900 channels in the C+L-band, compared to a mere five CWDM channels in that same range.
@@
An even more flexible form of DWDM has emerged in recent years—namely “elastic” or “flex-grid” optical networking [43, 44]. Rather than relying on fixed channel center frequencies and widths, flex-grid supports spectral channels of a variety of widths, enabling precise allocation of bandwidth to support disparate user needs; e.g., a high-bandwidth demand can receive a single ultrabroad 1 THz channel, while a low-bandwidth user concurrently obtains a 12.5 GHz one—there is no need to force them onto a fixed grid that may leave critical resources underutilized. ITU recommends central frequencies on a 6.25 GHz grid and slot widths in multiples of 12.5 GHz, but otherwise allows for any set of nonoverlapping channels [42]. In practice, flex-grid implementations are enabled by technology such as wavelength-selective switches (WSSs) that combine Fourier-transform pulse shaping [45, 46] with sub-spot-size gratings for fully programmable control of the amplitude, phase, and output fiber on a frequency-by-frequency basis (limited only by spectral resolution) [47, 48]. With such technology, it is possible to not only realize specific instantiations of flexible DWDM allocations, but also reconfigure slots on demand as bandwidth requirements evolve.
@@
An additional multiplexing approach available to optical networks, time-division multiplexing (TDM) assigns users slots in the time domain, typically interleaved such that $N$ users each communicating at rate $R_{0}$ multiplex into a single line rate of $NR_{0}$. Also enjoying a longstanding history in lightwave communications [49, 50, 51], TDM can be divided into electrical and optical versions, distinguished by whether the multiplexing and de-multiplexing operations are performed in the electronic or photonic domains, respectively. In the former case, multiple data streams are combined digitally and modulated onto a optical field; at the receiver, the data stream is detected at the full line rate and then demultiplexed into constituent tributary channels digitally. In the optical approach, the tributary channels are modulated onto optical fields separately, then combined, transmitted, and demultiplexed optically before optical-electrical (OE) conversion at separate detectors. As an entirely digital technique, electrical TDM requires no modifications to the optical architecture, yet the line rate $NR_{0}$ is then constrained by the bandwidth of OE conversion (either on the modulation or detection front). By contrast, optical TDM requires OE bandwidths only at the tributary rate $R_{0}$, thus in principle enabling much higher total line rates. Nevertheless, all-optical TDM still requires resources at the full bandwidth $NR_{0}$, including short pulses, delay lines, and fast switches. Such technological difficulties combined with the continued increase in digital signal processing (DSP) bandwidths have prevented optical TDM from making a significant commercial impact.
@@
Finally, moving beyond standard SMF to novel fiber types is opening up additional multiuser opportunities in SDM, where users are assigned distinct transverse spatial modes in, e.g., the same core of a multimode fiber (MMF) or separate cores in the common cladding of a multicore fiber [52, 8, 53]. While the cost of system upgrades makes the level of commercial SDM adoption unclear at the moment, SDM’s ability to multiply the transmission capacity of individual cables by potentially orders of magnitude suggests a compelling niche in space-constrained and high-bandwidth links.